Jaime is twice as old as his brother, Ricardo. In 5 years, Ricardo will be the same age as his brother was ten years ago. What are their present ages?
Thank you in advance for any help

Answer by jorel1380(3719)

j=2r
r+5=j-10
r=j-15
r=2r-15
r=15
j=30..